Equipment and Calibration History
Page: 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26 27 28 29 30 31 32 33 [34] 35 36 37 38 39 40 41 42 43 44 45 46 47 48 49 50 51 52 53 54 55 56 57 58 59 60 61 62 63 64 65 66 67 68 69 70Ser No. | Date | Loc From | Title | Loc To | Notes |
---|---|---|---|---|---|
LI-7500A 75H-2669 | 2019-01-31 | Sherman Barn Gas Analyzer CO2/H2O | Removed | Lab none | Removed from SB and brought back to the lab for calibration. It was replaced by 75H-2180. |
LI-7500A 1594 | 2019-01-31 | Lab none | Installed | Bouldin Corn Gas Analyzer CO2/H2O | Installed at BC to replace sn 75H-2668, which was brought back to the lab for calibration. |
LI-7500A 75H-2668 | 2019-01-31 | Bouldin Corn Gas Analyzer CO2/H2O | Removed | Lab none | Removed from BC and brought back to the lab for calibration. It was replaced by sn 1594. |
LI-7500A 0035 | 2019-01-31 | Lab none | Installed | Bouldin Alfalfa Gas Analyzer CO2/H2O | Installed at BA to replace sn 75H-2667, which was brought back to the lab for calibration. |
LI-7500A 75H-2667 | 2019-01-31 | Bouldin Alfalfa Gas Analyzer CO2/H2O | Removed | Lab none | Removed from BA and brought back to the lab for calibration. It was replaced by sn 0035. |
CR1000 50889 | 2019-01-31 | East Pond Data Logger | Serviced | East Pond Data Logger | Changed Pakbus address from 1 to 89. |
HMP60 L1740420 | 2019-01-31 | Lab none | Installed | East End RH & Temp | Installed at EE to replace HMP45 sn U3030046, which was removed for bad RH readings. |
HMP45AC U3030046 | 2019-01-31 | East End RH & Temp | Removed | Lab none | Removed from EE b/c RH was not working (ranged only b/t 10 and 13%). It was replaced by HMP60 sn L1740420. |
CR1000 47109 | 2019-01-31 | West Pond Data Logger Met | Serviced | West Pond Data Logger Met | Changed Pakbus address from 1 to 9. |
Mark I A000251 | 2017-09-07 | Lab none | Installed | Twitchell Alfalfa Weather Station | Installed at Twitchell Alfalfa with Joe Berry. This was just a quick stop on our way to Tonzi and Vaira. |
Line Quantum LQ1570 | 2019-01-29 | Factory | Disposed | Disposed | Disposed of at factory. Licor determined it was necessary to replace practically the whole innards of the sensor. |
Line Quantum LQ1570 | 2019-01-16 | Lab none | Shipped | Factory | Joe had cleaned up and tested this sensor, but we saw only noise in the data, so we returned to Licor for inspection and repair, RMA 211671. |
LI-7500A 1594 | 2019-01-30 | Lab none | Calibration | Lab none | Last calibration: 2018-10-03 AGC=98.88 SDM=0 DAC1=0 DAC2=0 Ambient P=100.07kPa, Tair=27.07C, Tdew=10.85C Zero Air at 1200psi, AMP Air 408.97 at 825 Tdewspan=16.00 CO2 zero 0.8644 -> .8643 Offset: -.16ppm CO2 span 1.015 -> 1.0157 Offset: .5ppm H2O zero 0.887 -> .8891 Offset: 11.9ppm H2O span 1.0056 -> 1.0043 Offset: .02ppm Notes: Calibrated by Jiangong. |
LI-7500A 0035 | 2019-01-28 | Lab none | Calibration | Lab none | Last calibration: 2018-10-02 AGC=99.2 SDM=0 DAC1=0 DAC2=0 Ambient P=100.7kPa, Tair=24.0C, Tdew=10.04C Zero Air at 1200psi, AMP Air 408.97 at 850 Tdewspan=16.00 CO2 zero 0.9375 -> .9375 Offset: .41ppm CO2 span 1.0142 -> 1.0193 Offset: 2.5ppm H2O zero 0.8981 -> .8969 Offset: -7.6ppm H2O span 1.0277 -> 1.0023 Offset: .54ppm Notes: Calibrated by Jiangong. Recorded on the from for sn 0035 but the correct config file (sn 0065) was loaded before the calibration, so the calibration is still good. |
LI-7500A 75H-2180 | 2019-01-25 | Lab none | Calibration | Lab none | Last calibration: 2018-10-02 AGC=98 SDM=0 DAC1=0 DAC2=0 Ambient P=101.2kPa, Tair=26.9C, Tdew=5.3C Zero Air at 1200psi, AMP Air 408.97 at 875 Tdewspan=16.00 CO2 zero 0.9137 -> .9137 Offset: 407.3ppm CO2 span 1.023 -> 1.0262 Offset: -1.67ppm H2O zero 0.883 -> .8891 Offset: 38.5ppm H2O span 0.9893 -> 1.0199 Offset: -.63ppm Notes: Date of last zeros were 2018-06-25, but the existing coefficients matched database entry from 2018-10-02. Training Jiangong. |
CR1000 47108 | 2019-01-24 | Sherman Barn Data Logger Met | Serviced | Sherman Barn Data Logger Met | Changed Pakbus address from 1 to 8. |
CR1000 38211 | 2019-01-24 | Sherman Wetland Data Logger Met | Serviced | Sherman Wetland Data Logger Met | Changed Pakbus address from 1 to 11. |
HMP45AC W0730036 | 2019-01-24 | Lab none | Installed | Sherman Wetland RH & Temp | Installed at SW to replace T2130005, which was brought back to the lab for inspection. |
HMP45AC T2130005 | 2019-01-24 | Sherman Wetland RH & Temp | Removed | Lab none | Brought back to the lab from SW for inspection b/c of data drops/spikes. It was replaced by sn W0730036. |
CR1000 47110 | 2019-01-24 | East End Data Logger Met | Serviced | East End Data Logger Met | Changed Pakbus address from 1 to 12. |
HMP45AC U3030046 | 2019-01-24 | Lab none | Installed | East End RH & Temp | Installed at EE to replace G4540023, which was brought back to the lab for inspection. |
HMP60 G4540023 | 2019-01-24 | East End RH & Temp | Removed | Lab none | Brought back to the lab from EE for inspection b/c of data drops/spikes. It was replaced by HMP45 sn U3030046. |
CR1000 38212 | 2019-01-24 | Mayberry Data Logger Tule | Serviced | Mayberry Data Logger Tule | Changed Pakbus address from 1 to 12. |
HMP45A U2120013 | 2019-01-24 | Mayberry RH & Temp | Serviced | Mayberry RH & Temp | Short in the 12V wire at the splice point. Repaired in the field with a screw terminal, but should be brought back to the lab for a full repair. |
HMP45AC W0730036 | 2019-01-23 | Lab none | Maintenance | Lab none | Spliced on a longer cable so total length is ~20'. New wiring colors are: Orange-Temperature, Blue-RH, White-SignalReference, Red-12V, Black-G, Bare/Shield-AG |
HMP45AC U3030046 | 2019-01-23 | Lab none | Maintenance | Lab none | Spliced on a longer cable so total length is ~24'. New wiring colors are: Orange-Temperature, Blue-RH, White-SignalReference, Green-Control, Red-12V, Black-G, Bare/Shield-AG |
CR10X 20138 | 2019-01-13 | Lab none | Installed | Tonzi | Installed near the Floor solar panels to record temperatures from a concrete cal panel for the FLIR thermal camera. |
LI-7500A 0073 | 2019-01-18 | Lab none | Installed | East Pond Gas Analyzer CO2/H2O | Installed at EP to replace sn 0035, which was brought back to the lab for calibration. |
LI-7500A 0035 | 2019-01-18 | East Pond Gas Analyzer CO2/H2O | Removed | Lab none | Removed from EP and brought back to the lab for calibration. it was replaced by sn 0073. |
LI-7500A 0418 | 2019-01-18 | Lab none | Installed | Sherman Wetland Gas Analyzer CO2/H2O | Installed at SW to replace 75H-2180, which was brought back to the lab for calibration. |
LI-7500A 75H-2180 | 2019-01-18 | Sherman Wetland Gas Analyzer CO2/H2O | Removed | Lab none | Removed from SW and brought back to the lab for calibration. It was replaced by sn 0418. |
LI-7500A 75H-2514 | 2019-01-18 | Lab none | Installed | Mayberry Gas Analyzer CO2/H2O | Installed at MB to replace sn 1594, which was brought back to the lab for calibration. |
LI-7500A 1594 | 2019-01-18 | Mayberry Gas Analyzer CO2/H2O | Removed | Lab Gas Analyzer CO2/H2O | Removed from MB and brought back to the lab for calibration. It was replaced by sn 75H-2514. |
LI-7500A 0073 | 2019-01-07 | Lab none | Calibration | Lab none | Last calibration: 2018-09-26 AGC=100 SDM=0 DAC1=0 DAC2=0 Ambient P=100.9kPa, Tair=18.7C, Tdew=9.5C Zero Air at 1200psi, AMP Air 408.97 at 850 Tdewspan=16.00 CO2 zero 0.9569 -> .9560 Offset: -2.67ppm CO2 span 1.0112 -> 1.0081 Offset: 1.65ppm H2O zero 0.7314 -> .7316 Offset: .6ppm H2O span 1.0183 -> 1.0023 Offset: .35ppm Notes: |
LI-7500A 0418 | 2019-01-07 | Lab none | Calibration | Lab none | Last calibration: 2018-10-10 AGC=97.2 SDM=0 DAC1=0 DAC2=0 Ambient P=100.9kPa, Tair=17.9C, Tdew=9.7C Zero Air at 1200psi, AMP Air 408.97 at 900 Tdewspan=16.00 CO2 zero 0.8982 -> .8992 Offset: 3.69ppm CO2 span 0.9965 -> 1.0033 Offset: -3.58ppm H2O zero 0.8454 -> .8465 Offset: 6.1ppm H2O span 0.9813 -> .9832 Offset: -.03ppm Notes: |
LI-7500A 75H-2514 | 2019-01-07 | Lab none | Calibration | Lab none | Last calibration: 2018-09-26 AGC=102.5 SDM=0 DAC1=0 DAC2=0 Ambient P=100.9kPa, Tair=18.0C, Tdew=10.1C Zero Air at 1200psi, AMP Air 408.97 at 850 Tdewspan=16.00 CO2 zero 0.9045 -> .9051 Offset: 2.36ppm CO2 span 1.0036 -> 1.0167 Offset: -6.83ppm H2O zero 0.837 -> .8397 Offset: 19.5ppm H2O span 1.0165 -> 1.0177 Offset: -.03ppm Notes: |
LI-7700 TG1-0421 | 2018-11-07 | Lab none | Installed | Bouldin Alfalfa Gas Analyzer CH4 | Installed at BA. The previous LI-7700 sn TG1-0424 was removed for a power board swap. |
LI-7500A 0041 | 2019-01-04 | Lab none | Installed | Tonzi-Floor | Installed on the Tonzi Floor to replace sn 75H-2514 for calibration. |
LI-7500A 75H-2514 | 2019-01-04 | Tonzi-Floor Gas Analyzer CO2/H2O | Removed | Lab none | Removed from the Tonzi Floor for calibration. Replaced by sn 0041 |
LI-7500A 0042 | 2019-01-04 | Lab none | Installed | Tonzi-Tower Gas Analyzer CO2/H2O | Installed on the Tonzi Tower to replace sn 0073 for calibration. |
LI-7500A 0073 | 2019-01-04 | Tonzi-Tower Gas Analyzer CO2/H2O | Removed | Lab Gas Analyzer CO2/H2O | Removed from the Tonzi Tower for calibration. Replaced by sn 0042 |
RF400 21519 | 2019-01-04 | Vaira | Removed | Lab none | Remove from the pond data logger box for testing in the lab. The antenna cable was broken. The radio has not been used in the field for a long while. |
RF400 21520 | 2019-01-04 | Vaira | Removed | Lab none | Remove from the met data logger box for testing in the lab. The radio has not been used in the field for a long while. |
LI-7500A 75H-2176 | 2019-01-04 | Lab none | Installed | Vaira Gas Analyzer CO2/H2O | Installed on the Vaira tower to replace sn 0418 for calibration. |
LI-7500A 0418 | 2019-01-04 | Vaira Gas Analyzer CO2/H2O | Removed | Lab none | Removed from the Vaira tower for calibration. Replaced by sn 75H-2176. |
GP-001 GP20140019 | 2019-01-03 | East End CO2 probe | Removed | Lab none | Removed and brought back to the lab for testing. |
LI-7500A 0041 | 2019-01-02 | Lab none | Calibration | Lab none | Last calibration: 2018-07-03 AGC=98 SDM=0 DAC1=0 DAC2=0 Ambient P=101.2kPa, Tair=20.1C, Tdew=3.7C Zero Air at 1400psi, AMP Air 408.97 at 875 Tdewspan=16.00 CO2 zero 0.9966 -> .9953 Offset: -4.19ppm CO2 span 1.0173 -> 1.0194 Offset: -1.25ppm H2O zero 0.8192 -> .8248 Offset: 34.0ppm H2O span 1.0285 -> 1.0312 Offset: -.07ppm Notes: H2O zero offset larger than I expected, but H2O span seems fine. |
LI-7500A 0042 | 2019-01-02 | Lab none | Calibration | Lab none | Last calibration: 2018-07-10 AGC=98.7 SDM=0 DAC1=0 DAC2=0 Ambient P=101.3kPa, Tair=19.2C, Tdew=.7C Zero Air at 1400psi, AMP Air 408.97 at 875 Tdewspan=16.00 CO2 zero 0.9485 -> .9489 Offset: 1.45ppm CO2 span 1.0039 -> 1.0187 Offset: -7.39ppm H2O zero 0.8406 -> .8434 Offset: 18.2ppm H2O span 1.0158 -> 1.0334 Offset: -.37ppm Notes: Dewpoint generator was set at 16.03C and I adjusted it to 16.00 before beginning calibration. H2O zero offset larger than I expected, but H2O span seems fine. |
LI-7500A 75H-2176 | 2019-01-02 | Lab none | Calibration | Lab none | Last calibration: 2018-07-10
AGC=102 SDM=0 DAC1=0 DAC2=0 Ambient P=101.2kPa, Tair=20.5C, Tdew=3.3C Zero Air at 1250psi, AMP Air 408.97 at 850 Tdewspan=16.00 CO2 zero 0.8671 -> .8676 Offset: 2.45ppm CO2 span 1.0029 -> 1.0114 Offset: -4.48ppm H2O zero 0.848 -> .8525 Offset: 27.5ppm H2O span 0.9088 -> .8206 Offset: -.27ppm Notes: H2O zero offset larger than I expected, but H2O span seems fine. |
BB-Fab BB-TZ_soilco2 | 2018-12-13 | Tonzi | Upgrade | Tonzi | This system has been failing overnight partly because one set of Vaisala probes have not been cycling off. I also added to old batter most recently used on East End. |
BB-Fab BB-TF | 2018-12-13 | Tonzi-Floor | Upgrade | Tonzi-Floor | Nine old G27 flooded batteries, a mixture of ages, were removed. Most tested bad. Two tested ok at first but weak later. They were replaced by four G31 flooded batteries wired in pairs for a 24V system. The panels and tower ends were also modified for 24V. |
RF401 35785 | 2018-12-19 | Tonzi | Removed | Lab none | Removed from the OSU SP1 system to test out data collection by LattePanda in the lab. |
RF401 35754 | 2018-12-19 | Tonzi | Removed | Lab none | Removed from the OSU SP3 system to test out data collection by LattePanda in the lab. |
Mark I A000251 | 2018-12-18 | Lab none | Installed | Bouldin Alfalfa Weather Station A | The Arable Mark was reinstalled after charging in the lab. |
Mark I A000251 | 2018-12-13 | Bouldin Alfalfa Weather Station A | Removed | Lab none | Brought back to the lab because it did not seem to be charging in the field. Seemed to charge up fine in the lab. |
SI-121 1282 | 2010-12-09 | Tonzi | Removed | Lab none | Removed from Tonzi with the end of the surface temperature experiment. |
SI-121 1281 | 2010-12-09 | Tonzi | Removed | Lab none | Removed from Tonzi with the end of the surface temperature experiment. |
LI-7500A 0041 | 2002-02-08 | Lab none | Installed | Tonzi-Floor | Installed on the Tonzi Tower to replace sn 0035 for calibration |
LI-7500A 0035 | 2002-02-08 | Tonzi-Floor | Removed | Lab none | Removed from the Tonzi Floor for calibration. Replaced by sn 0041. |
NR-LITE 990350 | 2002-02-08 | Lab none | Installed | Tonzi Gas Analyzer CO2/H2O | Installed inside the orange fence on the floor met CR10x data logger. |
LI-7500A 0073 | 2002-02-08 | Lab Gas Analyzer CO2/H2O | Installed | Tonzi-Tower | Installed on the Tonzi Tower to replace sn 0042 for calibration |
LI-7500A 0042 | 2002-02-08 | Tonzi-Tower Gas Analyzer CO2/H2O | Removed | Lab Gas Analyzer CO2/H2O | Removed from the Tonzi Tower for calibration. Replaced by sn 0073. |
DO6400/T DO-06 | 2018-12-06 | East End DO Sensor | Removed | Lab none | Removed from East End because it was not working and has been made obsolete by MiniDot sensor. |
Mark I A000251 | 2018-11-28 | Lab none | Installed | Bouldin Alfalfa Weather Station A | Installed at BA to replace sn A000591, which will be sent back to Arable for inspection. |
Mark I A000591 | 2018-11-28 | Bouldin Alfalfa Weather Station A | Removed | Lab none | Removed from BA and brought back to the lab. We will send this back to Arable b/c they noticed a faulty thermocouple. It was replaced by A000251. |
HMP60 L1740420 | 2018-11-14 | Factory | Received | Lab none | Received back from Campbell Sci. They tested it against their reference for temp and RH at 90%, 50%, and 20% and could not find any issues. We should test it again in the lab. |
HMP60 L1740398 | 2018-11-14 | Factory | Received | Lab none | Received back from Campbell Sci. They tested it against their reference for temp and RH at 90%, 50%, and 20% and could not find any issues. We should test it again in the lab. |
HMP60 L2310736 | 2018-09-14 | Lab none | Disposed | Factory | Shipped back to Campbell Sci for testing (low readings). They found the sensor to be irreparably bad and disposed of it for us. |
HMP60 L1740420 | 2018-09-14 | Lab none | Shipped | Factory | Shipped back to Campbell Sci for testing (drifting downward). |
HMP60 L1740398 | 2018-09-14 | Lab none | Shipped | Factory | Shipped back to Campbell Sci for testing (drifting downward). |
BB-Fab BB-TF | 2018-10-29 | Tonzi-Floor | Maintenance | Tonzi-Floor | One of the batteries in the battery bank seems to have an internal short and was sapping power. The bad battery was removed. Power started to be bad on October 20. |
GP-001 GP20140016 | 2018-11-06 | Lab none | Tested | Lab none | Connected to a CR6 and recorded 10s data. The sensor was maxing out (reading ~5V) on both 0-2% and 0-6% ranges. The connection between sensor and cable seemed secure and clean. The sensor should go back to the factory. |
GP-001 GP20140018 | 2018-11-06 | Lab none | Tested | Lab none | Connected to a CR6 and recorded 10s data. The sensor was maxing out (reading ~5V) on both 0-2% and 0-6% ranges. The connection between sensor and cable seemed secure and clean. The sensor should go back to the factory. |
LI-7700 TG1-0421 | 2018-11-06 | Lab none | Calibration | Lab none | Calibration Location: Lab Last calibration: 2018-05-14 RSSI=81%, Optical RH=13% Ambient P=100.2kPa, Tair=26.7C Zero Air at 1350psi, AMP Air 1.9497 at 950 CH4 Zero Offset: -0.015ppm CH4 Span Offset: 0.0083ppm Notes: First calibration after power board replacement |
LI-7700 TG1-0421 | 2018-11-05 | Lab none | Upgrade | Lab none | I swapped out power board pn 9677-079G per Licor's instructions. The new power board is pn 9977-079. According to Licor, "the original board needs a new resistor for the "under-voltage" problem (potential to drop out when heater and washers run simultaneously while incoming power is lower than usual, say 11-10 VDC)". The new board Rev 0A (V3) should not have that problem. |
GP-001 GP20140026 | 2018-11-01 | Lab none | Installed | West Pond CO2 Probe B | Installed at WP to replace sn GP20140018, which was brought back to the lab for too-high readings. |
GP-001 GP20140018 | 2018-10-25 | West Pond CO2 Probe B | Removed | Lab none | Removed from WP because of high readings. It was replaced by sn GP20140026 two weeks later. |
LI-7700 TG1-0423 | 2018-11-01 | Lab none | Installed | East End Gas Analyzer CH4 | Installed at EE to replace sn TG1-0421, which was brought back to the lab for a power board replacement and calibration. |
LI-7700 TG1-0421 | 2018-11-01 | East End Gas Analyzer CH4 | Removed | Lab none | Removed from EE and brought back to the lab for a power board replacement and calibration. It was replaced by sn TG1-0423 |
GP-001 GP20140084 | 2018-11-01 | Lab none | Installed | Sherman Wetland CO2 Probe | Installed at SW to replace sn GP20140016, which was removed b/c it was maxing out. |
Mark I A000591 | 2018-11-02 | Bouldin Alfalfa Weather Station A | Notes | Bouldin Alfalfa Weather Station A | Arable e-mailed to say sn A000591 which has a bad backup temperature sensor and requested we send it back. Not knowing what the exact source of the failure is, and having checked and found no similar behavior in any other device, they want to find out why. We planned to remove this Arable on 2018-10-04, but instead removed sn A000251 from BC and sent that Arable back to the factory. The Arable at BC was replaced by sn A001911. sn A000591 is still at BA for now. |
Mark I A000251 | 2018-10-04 | Bouldin Corn Weather Station A | Removed | Lab none | Removed from Bouldin Corn to be sent back to Arable for inspection of back up temperature sensor. Replaced by sn A001911. Later we realized we removed the incorrect Arable and they had requested instead sn A000591. |
LI-7700 TG1-0423 | 2018-10-30 | Lab none | Calibration | Lab none | Calibration Location: Lab Last calibration: 2018-05-08 RSSI=80%, Optical RH=8% Ambient P=100.8kPa, Tair=27.97C Zero Air at 1500psi, AMP Air 1.9497 at 1000 CH4 Zero Offset: -0.015ppm CH4 Span Offset: -0.0027ppm Notes: First calibration after power board was replaced |
LI-7700 TG1-0423 | 2018-10-29 | Lab none | Upgrade | Lab none | I swapped out power board pn 9677-073G per Licor's instructions. The new power board is pn 9977-079. According to Licor the original boards Rev 0, could be damaged by exposure to voltage above 24VDC. The new board Rev 0A (V3) should not have that problem.
Note: This instrument was replaced with 2 original screws and 2 long screws. The long screws were used to help re-assemble the instrument, and I forgot to swap them for the original, shorter, screws. |
LI-7500A 75H-2150 | 2018-10-28 | Lab none | Installed | East End Gas Analyzer CO2/H2O | Installed at EE to replace sn 75H-2176, which was brought back to the lab for calibration. |
LI-7500A 75H-2176 | 2018-10-28 | East End Gas Analyzer CO2/H2O | Removed | Lab none | Removed from EE and brought back to the lab for calibration. It was replaced by sn 75H-2150. |
NR-LITE 990354 | 2018-10-30 | Lab none | Location | Lab none | This sensor is on the tram and has probably been there the last 10 years. |
PAR-LITE 030381 | 2018-10-30 | Lab none | Location | Lab none | This sensor has been on the tram in the outgoing position for at least the last 10 years. |
PAR-LITE 030380 | 2018-10-30 | Lab none | Location | Lab none | This sensor has been on the tram in the incoming position for at least the last 10 years. |
LI-7700 TG1-0424 | 2018-10-25 | Lab none | Installed | Sherman Wetland Gas Analyzer CH4 | Installed at SW to replace sn TG1-0423, which was brought back to the lab for power board replacement and for calibration. Heater settings changed from Bouldin settings (top: 25%, 2 degC, 18:00-11:00; bottom: 25%, 30%, 18:00-11:00) to Sherman/Twitchell settings (top: 5%, 2 degC, 18:00-9:00; bottom: 5%, 30%, 18:00-9:00). |
LI-7700 TG1-0423 | 2018-10-25 | Sherman Wetland Gas Analyzer CH4 | Removed | Lab Gas Analyzer CH4 | Removed from SW and brought back to the lab for a power board swap and calibration. It was replaced by sn TG1-0424. |
LI-7500A 75H-2182 | 2018-10-25 | Lab none | Installed | West Pond Gas Analyzer CO2/H2O | Installed at WP to replace sn 0042, which was brought back to the lab for calibration. |
LI-7500A 0042 | 2018-10-25 | West Pond Gas Analyzer CO2/H2O | Removed | Lab none | Removed from WP and brought back to the lab for calibration. it was replaced by sn 75H-2182. |
TR-5251 40415-1006 | 2018-10-25 | Mayberry Precipitation | Tested | Mayberry Precipitation | Repaired chewed cable and tested with 475ml water. Got 100 tips. Did not adjust set screws. |
PAR-LITE 040412 | 2018-10-25 | Lab none | Installed | Mayberry PAR Sensor Reflected Water | Installed at MB (measuring reflected water) to replace sn 010161, which was brought back to the lab for repair/calibration. |
PAR-LITE 010161 | 2018-10-25 | Mayberry PAR Sensor Reflected Water | Removed | Lab none | Removed from MB (measuring reflected water) and brought back to the lab to replace the rotted cable and for calibration. It was replaced by sn 040412. |
PAR-LITE 010159 | 2018-10-25 | Lab none | Installed | Mayberry PAR Sensor Incoming | Installed at MB in the incoming position to replace sn 10162, which was brought back to the lab for repair/calibration. |
PAR-LITE 10162 | 2018-10-25 | Mayberry PAR Sensor Incoming | Removed | Lab none | Removed from MB and brought back to the lab to replace the rotted cable and for calibration. It was replaced by sn 010159. |
LI-7700 TG1-0211 | 2018-10-25 | East Pond Gas Analyzer CH4 | Calibration | East Pond Gas Analyzer CH4 | Calibration Location: East Pond
Last calibration: 2018-04-13 RSSI=74%, Optical RH=10% Ambient P=101.6kPa, Tair=26.4C Zero Air at 1950psi, AMP Air 3.121 at 2100 CH4 Zero Offset: .029ppm CH4 Span Offset: .006ppm Notes: Calm day |
LI-7700 TG1-0155 | 2018-10-25 | Bouldin Corn Gas Analyzer CH4 | Calibration | Bouldin Corn Gas Analyzer CH4 | Calibration Location: Bouldin Corn
Last calibration: 2018-04-04 RSSI=91%, Optical RH=16% Ambient P=101.5kPa, Tair=27.2C Zero Air at 1925psi, AMP Air 3.121 at 1975 CH4 Zero Offset: .015ppm CH4 Span Offset: -.013ppm Notes: Gently breezy day |